# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

### v2

@@ I Introduction
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the Kullback-Leibler (KL) divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product.
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
The proof of this lemma can be found in Appendix -A. This lemma states that the stationary points for maximizing the expected Gaussian likelihood are achieved when the Gaussian distribution matches the mean and variance of the given distribution. Therefore, as a special case of (10), solving (9a) requires to match the first and second order moment:
@@
$+\mathbb{D}\left\{\xi_{t-1}\right\}-A\hat{x}_{t-1|t-1}\hat{x}_{t-1|t-1}^{\top}A^{\top},$
@@ III-C Structure of Stationary Points of $J(\hat{x}_{t},P_{t})$
$+\frac{1}{2}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)^{\mathrm{T}}P_{t|t-1}^{-1}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)$
@@
$=$ $\Sigma^{-1}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{(x-\mu)(x-\mu)^{\top}f(x)\right\}$
@@ IV Natural Gradient Gasussian Approximation
$=$ $\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\big{\{}(x_{t}-\hat{x}_{t}^{(i)})(x_{t}-\hat{x}_{t}^{(i)})^{\top}\ell(x_{t},y_{t})\big{\}}$
@@
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $P_{t|t-1}^{-1}+\left(P_{t}^{-1}\right)^{(i)}\cdot\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\left(x_{t}-\hat{x}_{t}^{(i)}\right)\left(x_{t}-\hat{x}_{t}^{(i)}\right)^{\top}\ell(x_{t},y_{t})\right\}-\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\ell(x_{t},y_{t})\right\},$ (28)
@@ V-A Convergence Analysis
This theorem indicates that natural gradient descent iteration in the update step of the NANO filter provides a guarantee of local convergence. This guarantee is achieved by approximating the objective function with second-order accuracy around the mean and first-order accuracy around the inverse of the covariance matrix.
The proof of Theorem 1 is provided in Appendix -B. The key idea is to show that the difference in update costs between consecutive iterations can be expressed as a semi-negative definite quadratic form. Technically, achieving this semi-negative definiteness relies on the use of the Fisher information matrix, which corrects the gradient direction to provide the steepest descent on the Gaussian manifold. This adjustment makes the gradient “natural” because it aligns with the geometry of the Riemannian space of Gaussian parameters.
@@
The proof of this corollary is provided in Appendix -C. Corollary 1 holds regardless of the initialization of the Gaussian parameters $\hat{x}^{(0)}_{t}$ and $P^{(0)}_{t}$, supporting the fact that the natural gradient is the steepest descent direction in the Gaussian manifold.
@@ V-B Stability Analysis
Next, we will analyze the stability of the proposed NANO filter. Stability is the most critical property, as it ensures that the estimation error remains bounded throughout the filtering process. Our stability analysis is conducted based on the state-space model given in (1). We focus on the case where the process and measurement noise are both zero-mean Gaussian noise, satisfying $\mathbf{\xi}_{t}\sim\mathcal{N}(0,Q_{t})$ and $\mathbf{\zeta}_{t}\sim\mathcal{\mathcal{missing}}{N}(0,R_{t})$. In this case, the measurement-dependent loss $\ell(x_{t},y_{t})$ is the log-likelihood loss function that satisfies $\ell(x_{t},y_{t})=-\log p(y_{t}|x_{t})=C_{0}+\frac{1}{2}(y_{t}-g(x_{t}))^{\top}R_{t}^{-1}(y_{t}-g(x_{t}))$. Here, $C_{0}>0$ is a constant value irrelevant to the state and measurement. For the purpose of stability analysis, we make the following regularity assumptions on the system functions $f$ and $g$.
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment. All the simulations and experiment are evaluated by root mean square error (RMSE) which is defined by
@@ VIII-B Nonlinear System Simulation: Air-traffic Control
$y_{t}$ $=\begin{bmatrix}\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}\\ \operatorname{atan}\left(\frac{p_{y,t}}{p_{x,t}}\right)\\ \operatorname{atan}\left(\frac{h}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}}}\right)\\ \frac{p_{x,t}\dot{p}_{x,t}+p_{y,t}\dot{p}_{y,t}}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}}\end{bmatrix}+\zeta_{t}.$
@@
Here, $q_{1}=2.5$ and $q_{2}=10^{-6}$. Similar to the linear case, we validate our method under two scenarios for the nonlinear system. In the first scenario, shown in Fig. 3, the NANO filter is compared with EKF, UKF, IEKF, and PLF. The results demonstrate that NANO filter significantly outperforms these established filters, achieving nearly a half reduction in RMSE. This substantial improvement highlights the NANO filter’s superior effectiveness in handling nonlinear dynamics. In the second scenario, we introduce measurement outliers to reflect conditions where the measurement data deviates from the assumed model. Specifically, we assume the actual measurement noise follows a mixture model: $\zeta_{t}\sim 0.9\cdot\mathcal{N}(\zeta_{t};0,R)+0.1\cdot\mathcal{N}(\zeta_{t};0,100\cdot R)$, indicating a 10% probability that measurements are contaminated by noise with a tenfold increase in standard deviation. As shown in Fig. 4, robust variants of the NANO filter—including the $\beta$-NANO, Huber-NANO, and Weight-NANO filters—significantly improve estimation accuracy in the presence of measurement outliers. Notably, the $\beta$-NANO filter with $\beta=10^{-2}$ exhibits the best performance, effectively mitigating the impact of outliers.
@@ IX Conclusion
-A Proof of Lemma 1
Proof.
$\log\mathcal{N}(x;\mu,\Sigma)=$ $-\frac{1}{2}\big{[}\log|\Sigma|$
$+(x-\mu)^{\top}\Sigma^{-1}(x-\mu)+\log(2\pi)^{n}\big{]},$
Combining these results, we have
Setting this derivative equal to zero gives
which implies (11b). ∎
-B Proof of Theorem 1
Proof.
$=$ $\left(P_{t}^{-1}\right)^{(i+1)}.$
According to (25), we have
$J_{t}^{(i+1)}-J_{t}^{(i)}$
$\leq$ $0$
-C Proof of Corollary 1
Proof.
$=$ $P_{t|t-1}^{-1}+C^{\top}R^{-1}C.$
$-P_{t}^{(1)}P_{t|t-1}^{-1}\left(\hat{x}_{t}^{(0)}-\hat{x}_{t|t-1}\right)$
$=$ $\hat{x}^{(0)}_{t}+P_{t|t}C^{\top}R^{-1}(y_{t}-C\hat{x}_{t}^{(0)})$
$=$ $\hat{x}_{t|t-1}+P_{t|t}C^{\top}R^{-1}\left(y_{t}-C\hat{x}_{t|t-1}\right).$
Furthermore, noting that:
$P_{t|t}C^{\top}R^{-1}$
$=$ $P_{t|t-1}C^{\top}R^{-1}$
$-P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1}CP_{t|t-1}C^{\top}R^{-1}$
$+(R+CP_{t|t-1}C^{\top})^{-1}R\big{)}R^{-1}$
$=$ $P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1},$
$\hat{x}^{(1)}_{t}=\hat{x}_{t|t-1}+K_{t}\left(y_{t}-C\hat{x}_{t|t-1}\right),$

### v3

@@ I Introduction
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the Kullback-Leibler (KL) divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product. The notation $\operatorname*{\mathrm{vec}}(A)$ refers to the vector obtained by stacking the columns of matrix $A$ into a single column vector. The trace of matrix $A$ is denoted as $\mathrm{Tr}(A)$. The notation $\mathbb{I}_{n\times n}$ represents the $n$-dimensional identity matrix.
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
Proof.
$\log\mathcal{N}(x;\mu,\Sigma)=$ $-\frac{1}{2}\big{[}\log|\Sigma|$
$+(x-\mu)^{\top}\Sigma^{-1}(x-\mu)+\log(2\pi)^{n}\big{]},$
Combining these results, we have
Setting this derivative equal to zero gives
which implies (11b). ∎
This lemma states that the stationary points for maximizing the expected Gaussian likelihood are achieved when the Gaussian distribution matches the mean and variance of the given distribution. Therefore, as a special case of (10), solving (9a) requires to match the first and second order moment:
@@
$+\mathbb{D}\left\{\xi_{t-1}\right\}-A\hat{x}_{t-1|t-1}\hat{x}_{t-1|t-1}^{\top}A^{\top}$
@@ III-C Structure of Stationary Points of $J(\hat{x}_{t},P_{t})$
$+\frac{1}{2}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)^{\top}P_{t|t-1}^{-1}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)$
@@
$=$ $\Sigma^{-1}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{(x-\mu)(x-\mu)^{\top}f(x)\right\}\Sigma^{-1}$
@@ IV Natural Gradient Gasussian Approximation
$=$ $\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\big{\{}(x_{t}-\hat{x}_{t}^{(i)})(x_{t}-\hat{x}_{t}^{(i)})^{\top}\ell(x_{t},y_{t})\big{\}}\left(P_{t}^{-1}\right)^{(i)}$
@@
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $P_{t|t-1}^{-1}+\left(P_{t}^{-1}\right)^{(i)}\cdot\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\Big{\{}(x_{t}-\hat{x}_{t}^{(i)})(x_{t}-\hat{x}_{t}^{(i)})^{\top}\ell(x_{t},y_{t})\Big{\}}\cdot\left(P_{t}^{-1}\right)^{(i)}-\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\ell(x_{t},y_{t})\right\},$ (28)
@@ V-A Convergence Analysis
Proof.
$=$ $\left(P_{t}^{-1}\right)^{(i+1)}.$
According to (25), we have
$J_{t}^{(i+1)}-J_{t}^{(i)}$
$\leq$ $0$
This theorem indicates that natural gradient descent iteration in the update step of the NANO filter provides a guarantee of local convergence. This guarantee is achieved by approximating the objective function with second-order accuracy around the mean and first-order accuracy around the inverse of the covariance matrix. The key idea of proof is to show that the difference in update costs between consecutive iterations can be expressed as a semi-negative definite quadratic form. Technically, achieving this semi-negative definiteness relies on the use of the Fisher information matrix, which corrects the gradient direction to provide the steepest descent on the Gaussian manifold. This adjustment makes the gradient “natural” because it aligns with the geometry of the Riemannian space of Gaussian parameters.
@@
Proof.
$=$ $P_{t|t-1}^{-1}+C^{\top}R^{-1}C.$
$-P_{t}^{(1)}P_{t|t-1}^{-1}\left(\hat{x}_{t}^{(0)}-\hat{x}_{t|t-1}\right)$
$=$ $\hat{x}^{(0)}_{t}+P_{t|t}C^{\top}R^{-1}(y_{t}-C\hat{x}_{t}^{(0)})$
$=$ $\hat{x}_{t|t-1}+P_{t|t}C^{\top}R^{-1}\left(y_{t}-C\hat{x}_{t|t-1}\right).$
Furthermore, noting that:
$P_{t|t}C^{\top}R^{-1}$
$=$ $P_{t|t-1}C^{\top}R^{-1}$
$-P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1}CP_{t|t-1}C^{\top}R^{-1}$
$+(R+CP_{t|t-1}C^{\top})^{-1}R\big{)}R^{-1}$
$=$ $P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1},$
$\hat{x}^{(1)}_{t}=\hat{x}_{t|t-1}+K_{t}\left(y_{t}-C\hat{x}_{t|t-1}\right),$
Corollary 1 holds regardless of the initialization of the Gaussian parameters $\hat{x}^{(0)}_{t}$ and $P^{(0)}_{t}$, supporting the fact that the natural gradient is the steepest descent direction in the Gaussian manifold.
@@ V-B Stability Analysis
Next, we will analyze the stability of the proposed NANO filter. Stability is the most critical property, as it ensures that the estimation error remains bounded throughout the filtering process. Our stability analysis is conducted based on the state-space model given in (1). We focus on the case where the process and measurement noise are both zero-mean Gaussian noise, satisfying $\mathbf{\xi}_{t}\sim\mathcal{N}(0,Q_{t})$ and $\mathbf{\zeta}_{t}\sim\mathcal{N}(0,R_{t})$. In this case, the measurement-dependent loss $\ell(x_{t},y_{t})$ is the log-likelihood loss function that satisfies $\ell(x_{t},y_{t})=-\log p(y_{t}|x_{t})=C_{0}+\frac{1}{2}(y_{t}-g(x_{t}))^{\top}R_{t}^{-1}(y_{t}-g(x_{t}))$. Here, $C_{0}>0$ is a constant value irrelevant to the state and measurement. For the purpose of stability analysis, we make the following regularity assumptions on the system functions $f$ and $g$.
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment [^1] The code of this paper can be found in https://github.com/TianyiMehdi/NANO-filter. . All the simulations and experiment are evaluated by root mean square error (RMSE) which is defined by
@@ VIII-B Nonlinear System Simulation: Air-traffic Control
$y_{t}$ $=\begin{bmatrix}\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}\\ \operatorname{atan}\left(\frac{p_{y,t}}{p_{x,t}}\right)\\ \operatorname{atan}\left(\frac{h}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}}}\right)\\ \frac{p_{x,t}v_{x,t}+p_{y,t}v_{y,t}}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}}\end{bmatrix}+\zeta_{t}.$
@@
Here, all noise parameters are the same as those in [14]. Similar to the linear case, we validate our method under two scenarios for the nonlinear system. In the first scenario, shown in Fig. 3, the NANO filter is compared with EKF, UKF, IEKF, and PLF. The results demonstrate that NANO filter significantly outperforms these established filters, achieving nearly a half reduction in RMSE. This substantial improvement highlights the NANO filter’s superior effectiveness in handling nonlinear dynamics. In the second scenario, we introduce measurement outliers to reflect conditions where the measurement data deviates from the assumed model. Specifically, we assume the actual measurement noise follows a mixture model: $\zeta_{t}\sim 0.9\cdot\mathcal{N}(\zeta_{t};0,R)+0.1\cdot\mathcal{N}(\zeta_{t};0,100\cdot R)$, indicating a 10% probability that measurements are contaminated by noise with a tenfold increase in standard deviation. As shown in Fig. 4, robust variants of the NANO filter—including the $\beta$-NANO, Huber-NANO, and Weight-NANO filters—significantly improve estimation accuracy in the presence of measurement outliers. Notably, the $\beta$-NANO filter with $\beta=10^{-2}$ exhibits the best performance, effectively mitigating the impact of outliers.